OpenStudy (anonymous):

Someone help me solve 3x-5=2x+6

OpenStudy (anonymous):

solving for x?

OpenStudy (anonymous):

yes

OpenStudy (nurali):

3x-5=2x+6 Subtract both side by 2x 3x-2x-5=2x-2x+6 x-5=6 add both side by 5 x-5+5=6+5 x=11

OpenStudy (dannyrod2000):

x=11

OpenStudy (nurali):

Welcome to Openstudy @AndreaEwen

OpenStudy (dannyrod2000):

i don't think she is new :P

OpenStudy (anonymous):

I am though haha i started like 2 hours ago at most 3

OpenStudy (dannyrod2000):

Welcome To Openstudy @cambam1323

OpenStudy (anonymous):

thanks haha

OpenStudy (dannyrod2000):

:)